OpenStudy (anonymous):

Find all zeros by factoring each function. f(x)= x^4 - x^2 - 30. Explain in detail please.

OpenStudy (mathmate):

Substitute y=x^2, so f(y)=y^2-y-30 After that, find x = sqrt(y) (note x could be complex)

OpenStudy (anonymous):

Complex?

OpenStudy (mathmate):

imaginary, when we take sqrt(-1), the result is i. Suppose we obtained y=4 or y=-3. Then \( x=\pm sqrt(4)\ or\ x=\pm sqrt(-3) \) which translates to \( x=-2,\ 2,\ -sqrt(3)i,\ or\ sqrt(3)i \)

OpenStudy (anonymous):

The final answer is sqrt(6), -sqrt(6), sqrt(5)i, -sqrt(5)i. but im trying to find out the steps my teacher took to find the answer.

OpenStudy (mathmate):

Start by factoring: g(y)=y^2-y-30

OpenStudy (anonymous):

Ok, I have that part, its (x^2 - 6)(x^2 + 5). Then i set the x's to zero, right? that's where i'm lost.

OpenStudy (mathmate):

Yes, y=x^2=6 => x^2=6, which can be solved as \( x=\pm\sqrt{6} \) You will similarly calculate the value of x for the other solution y=-5. (result is complex, see example above)

OpenStudy (mathmate):

Sorry, \( y=x^2-6=0 => x^2=6 => x=\pm\sqrt{6} \)

OpenStudy (anonymous):

So I factor again? Then is your complex answer an example or...?

OpenStudy (mathmate):

Similarly, \( y=x^2-(-5)=0 => x^2=-5 => x=\pm\sqrt{-5}=\pm \sqrt{5} i \) where \( i = \sqrt{-1} \)

OpenStudy (anonymous):

Ok, so that is the factored function? So => is a symbol for?

OpenStudy (mathmate):

=> means "implies", or "therefore"

OpenStudy (anonymous):

So if i wanted to write this on my paper i would write it as...

OpenStudy (anonymous):

But i'm still confused. After you factor the 1st part, you then factor what you factored?

OpenStudy (mathmate):

You can also write them on separate lines, one after another, as long as each line follows from the previous.

OpenStudy (mathmate):

Such as \( x^2=16\) \( x=\pm 2 \)

OpenStudy (anonymous):

Oh, Ok. Makes sense. But your answer is a little complicated. Can you break it down more please?

OpenStudy (mathmate):

where y+5=0?

OpenStudy (anonymous):

Yes, that problem was complicated.

OpenStudy (mathmate):

Sorry for being to condensed! \( (y+5)=0 \) \( (y=-5 \) \( x^2=-5 \) \( x=\pm \sqrt{-5} \) \( x=\pm \sqrt{5} \ i\) Do you follow the steps better now?

OpenStudy (mathmate):

If you don't, do let me know. It's no use to copy the step, get 100% for the homework, and get 10% for the exam!

OpenStudy (anonymous):

Alright, so the step you just put is after i factor the first time?

OpenStudy (mathmate):

Sorry, all the x's should have been y's. Deleting the previous post to avoid confusion. Yes. After the quadratic in y had been factored, there were two cases, either (y-6)=0, or (y+5)=0 This is for the second case, y+5=0. I believe you did the (y-6)=0 case.

OpenStudy (anonymous):

Yes, the (x-6) was what my teacher did. So let me break this down from the beginning, i factor the first part, then i factor again to solve for y? Then i get my initial answer?

OpenStudy (mathmate):

The final answer should consist of 4 values (for a quartic). Two of these correspond to (y-6)=0, and the other two from (y+5)=0.

OpenStudy (anonymous):

And your work above explains your answer right?

OpenStudy (mathmate):

Yes, you will need to combine the work for the two cases.

OpenStudy (anonymous):

Ok, Got you. So (x+5) = 0?

OpenStudy (anonymous):

Then x=-5, then x^2=-5, x= +- sqrt-5, x= +-sqrt5i? and then the same for x^2 -6=0?

OpenStudy (mathmate):

yep!